# Effect of freestream turbulence on the coherent dynamics of a wind turbine wake

### v1

@@ 7 Conclusion
Here $R_{21}^{{}^{\prime}}$ represents the streamwise autocorrelation function of transverse velocity fluctuation ($u_{2}^{\prime}$). $r\textbf{e}_{1}$ represents some displacement $r$ along the $\textbf{e}_{1}$ direction ($\textbf{e}_{1}$ is the unit vector along the streamwise direction). Finally $L_{v}$ is calculated by integrating $R_{21}^{{}^{\prime}}$ up to a certain distance $\hat{r}$ where $R_{21}^{{}^{\prime}}$ becomes zero. Measuring $\hat{r}$ is not trivial since the nature of $R_{21}^{{}^{\prime}}$ changes with streamwise distance. Figure 24 shows the variation of $R_{21}^{{}^{\prime}}$ along the the centreline ($y=0$, $z=0$), and $2D$ and $5D$ away from the grid SRG38. Closer to the grid, $R_{21}^{{}^{\prime}}$ clearly falls off to zero, while further downstream it does not. Therefore, a fitting function needs to be used. Note that the background correlation shows multiple secondary peaks, the strongest one is shown by the ‘+’ sign in figure 24(b). A cutoff twice the peak value was used (as shown by the red dashed line) and the data above the cutoff was used for fitting. A bi-Gaussian function was used for fitting (two separate Gaussian curves were fitted on both sides of the peak). As an example, the fitted curve is shown in figure figure 24(b) by the black line. Following equation 3, $L_{v}$ was calculated on both the sides and an average of the two was taken. Towards the ends of the FOV where fitting could not be performed on both sides, $L_{v}$ was estimated simply based on data from one side. Figure 25 shows the variation of integral length scale ($L_{v}$) along the geometrical centreline. Similar to figure 22, the filled contour shows transverse variation of $L_{v}$. The dashed vertical lines show the positions of the turbine in different experiments.
The integral length scales could also be obtained using other autocorrelation functions such as $R_{12}^{{}^{\prime}}$ (transverse autocorrelation function of streamwise velocity fluctuation) or $R_{11}^{{}^{\prime}}$ (streamwise autocorrelation function of streamwise velocity fluctuation). Examples of these autocorrelation functions downstream of the grids SRG38 and SFG5 are shown in figure 26. Looking at the area under the curves, one can tell that $R_{11}^{{}^{\prime}}$ yields much larger integral length scales while $R_{21}^{{}^{\prime}}$ and $R_{12}^{{}^{\prime}}$ yields integral length scales of similar size.
Figure 27: Example low frequency modes for the (a) no grid case and (b) case 4b that are not considered to be a part of the coherent component of the flow and are likely attributable to an intermittent large-scale phenomenon.
@@
Although mode clustering simplifies the OMD spectrum, care must be taken while selecting modes from the reduced spectrum (after performing mode clustering) to represent the coherent component of the flow. The lesser damped modes generally represent a physically meaningful mode, but drawing a cutoff on the growth rate ($Re(\lambda_{OMD})$) is subjective. Looking at the nature of the spectra, the modes $Re(\lambda_{OMD})\gtrsim-0.05$, or in that vicinity, were considered. Before selecting a mode, each mode shape had to be checked manually to see if they show any physically meaningful structure. Sometimes a slightly more damped mode had to be selected since its spatial coherence was more interpretable in a physically meaningful way. Occasionally some modes were not selected despite being less damped than some other selected modes (example figure 12(i)). The spatial shape of these modes revealed very large scale smeared structures which were likely due to the intermittent passage of large scale structures having a size smaller than that indicated by the mode shapes. Example of such modes are shown in figure 27 for the no grid case and case 4b. These modes are also highlighted by a blue arrow in the spectra shown in figures 12(c) and 12(i). Furthermore, when a mode was obtained by clustering other modes, the growth rate of the new mode was simply calculated as a weighted average of the growth rates of the modes forming the cluster. This was done simply to place the new mode somewhere in the spectrum and this growth rate need not have any physical meaning. In summary, the shape of the modes involved in clustering was always checked irrespective of their growth rates.

### v2

@@ 6 Conclusion
Once the clusters are identified, each of the modes in the cluster are stacked in a matrix (let us denote it as $M$) as column vectors multiplied by a significance parameter or weight $\sigma_{l}$. If the $k^{th}$ cluster has $n$ modes, then $M_{k}=[\sigma_{1}\phi_{1},\sigma_{2}\phi_{2}...\sigma_{n}\phi_{n}]$. Thereafter, a singular value decomposition of $M_{k}$ is performed and the first singular vector is used as the mode representing the cluster. However, it is worthwhile to note that several aspects of the method are still not strictly defined and deserve further consideration. One of them is how to assign a frequency and growth rate to a clustered mode. Further, which statistic is the best to use as $\sigma$? For example, one could project the snapshots on the OMD modes to obtain the time varying coefficients of the modes. The magnitude of the coefficient at a specific time or a norm of the coefficient could be used as $\sigma$. Here however, it was found that using such a statistic undermined the influence of the lesser damped modes within a cluster. Accordingly, a quantity related to the growth rate ($\Re(\lambda_{OMD})$) of the modes was thought to be a more reasonable measure of $\sigma$ motivating the use of the quantity $1/|\Re(\lambda_{OMD})|$. The dominant modes (growth rate closer to zero) are therefore assigned a higher significance value. This method showed faster convergence for the current data set and yielded better converged modes representing a cluster. The modes representing a cluster are also assigned a virtual growth rate and a frequency which is obtained by taking a weighted average of the frequencies and growth rates of the modes forming a cluster with $\sigma$ used as the weight. The whole process is repeated until no two modes are similar.
The choice of the cutoff parameters $\zeta_{spat}$ and $\zeta_{spec}$ also involves some subjectivity. The effects of $\zeta_{spat}$ and $\zeta_{spec}$ on the converged number of modes ($N_{c}$) for the four main cases are shown in figure 25. The rank of OMD was set to 175 (i.e. the total number of initial modes $N=175$) for all cases following our previous work [Biswas and Buxton, 2024b]. No mode reduction is achieved for $\zeta_{spat}\geq 0.96$, while for $\zeta_{spat}\leq 0.9$ the spectra are over simplified leaving too few modes. A different $\zeta_{spat}$ was selected for each case such that $N_{c}/N\approx 0.9$. For the cases shown in figure 25, the selected cutoffs are shown by the vertical dashed line. The different symbols in figure 25 correspond to different values of $\zeta_{spec}$. Note that the mode clustering process does not depend significantly on $\zeta_{spec}$ for $\zeta_{spec}\geq 0.2$. The analysis was therefore continued with $\zeta_{spec}=0.3$ for all the cases.
Figure 25: Variation of the number of converged modes ($N_{c}$) with $\zeta_{spat}$ for different $\zeta_{spec}$ for the cases (a) 1, (b) 2c, (c) 3b and (d) 4b.
Figure 26: Example low frequency modes for the (a) no grid case and (b) case 4b that are not considered to be a part of the coherent component of the flow and are likely attributable to an intermittent large-scale phenomenon. The real part of the OMD modes ($\Re(\phi_{OMD})$) is shown.
@@
Although mode clustering simplifies the OMD spectrum, care must be taken while selecting modes from the reduced spectrum (after performing mode clustering) to represent the coherent component of the flow. The lesser damped modes generally represent a physically meaningful mode, but drawing a cutoff on the growth rate ($\Re(\lambda_{OMD})$) is subjective. Looking at the nature of the spectra, the modes $\Re(\lambda_{OMD})\gtrsim-0.05$, or in that vicinity, were considered. Before selecting a mode, each mode shape had to be checked manually to see if they show any physically meaningful structure. Sometimes a slightly more damped mode had to be selected since its spatial coherence was more interpretable in a physically meaningful way. Occasionally some modes were not selected despite being less damped than some other selected modes (example figure 16(c)). The spatial shape of these modes revealed very large scale smeared structures which were likely due to the intermittent passage of large scale structures having a size smaller than that indicated by the mode shapes. Example of such modes are shown in figure 26 for the no grid case and case 4b. These modes are also highlighted by a blue arrow in the spectra shown in figures 16(c) and 16(l). Furthermore, when a mode was obtained by clustering other modes, the growth rate of the new mode was simply calculated as a weighted average of the growth rates of the modes forming the cluster. This was done simply to place the new mode somewhere in the spectrum and this growth rate need not have any physical meaning. In summary, the shape of the modes involved in clustering was always checked irrespective of their growth rates.